Connect Homework: Chapter 11 At September 30, the end of Beijing Companys third quarter, the following stockholders equity
Answer #1
Sep 30 Beg. Bal. Oct. 2 Oct. 25 Oct. 31 Nov. 5 Dec. 1 Dec. 31
Common stock $360000 $360000 $360000 $360000 (360000+40000)= $400000 $400000 $400000
Common stock dividend distributable 0 0 40000 0 0 0
Paid-in capital in excess of par, common stock 110000 110000 110000 (110000+43000)= 153000 153000 153000 153000
Retained earnings 360000 (360000-60000)= 300000 300000 (300000-83000)= 217000 217000 217000 (217000+260000)= 477000
Total $830000 $770000 $770000 $770000 $770000 $770000 $1030000